# Decoherence of electron spin qubit during transfer between two semiconductor quantum dots at low magnetic fields

### v2

@@ II.3 Dynamics in presence of coupling to environment
Electric field noise leads to fluctuations of mean position of electron localized in each of the dots. In presence of finite gradient of $z$ component of the magnetic field, these lead to fluctuations in Zeeman splittings $E_{Z,L(R)}$ in both QDs [28, 29]. Similar fluctuations are caused by hyperine interaction with nuclei in each of the dots. We will include these processes later in the analysis, and now we will focus on processes that are activated by interdot tranfer of the electron caused by detuning sweep. These requite the coupling of the environment to the charge states only:
@@
where the $\hat{V}_{\epsilon},\hat{V}_{t_{c}}$ operators acts on the environmental degrees of freedom and represent quantum fluctuations in detuning and tunnel coupling respectively. Such environment is typically assumed to be in thermal equilibrium at low, but finite electron temperature of $T\approx 100\,$mK such that $12t_{c}/k_{\text{B}}T>$ 1 [1]. The high-frequency environmental noise causes inelastic transitions between adiabatic states of Eq. (1) , which we model by Adiabatic Master Equation (AME) [36, 37, 38] of the form:
@@
where all of the operators are written in the adiabatic frame, ${\tilde{O}}=\hat{A}(t)\hat{O}(t)\hat{A}^{\dagger}(t)$, i.e. in the basis associated with the instantaneous states of $\hat{H}_{0}(t)$ from Eq. (1). By definition, the transformation associated with an operator $\hat{A}(t)$ diagonalizes the Hamiltonian at every time instant:
@@
Firstly, we consider spin-diagonal transitions between the orbital states. After transformation the operators ${\tilde{\tau}}_{s\pm}=|\tilde{\psi}_{s\pm}(t)\rangle\langle\tilde{\psi}_{s\mp}(t)|$ are responsible for such transitions between the eigenstate of energy ${E_{s\pm}=\sigma_{s}E_{Z}/2\pm\Omega_{s}/2}$ and the eigenstate with energy ${E_{s\mp}=\sigma_{s}E_{Z}/2\mp\Omega_{s}/2}$, respectively. With this process we associate corresponding time-dependent rates $\Gamma_{\pm,\text{charge}}^{(s)}(t)$ of excitation (relaxation) for spin-s. Secondly we add transitions between the eigenstates that are adiabatically linked to the opposite spins, i.e. ${\tilde{\sigma}}_{+}=|\tilde{\psi}_{\uparrow\pm}(t)\rangle\langle\tilde{\psi}_{\downarrow\pm}(t)|$, ${\tilde{\sigma}}_{-}=|\tilde{\psi}_{\downarrow\pm}(t)\rangle\langle\tilde{\psi}_{\uparrow\pm}(t)|$, with associated time-depedent transition rates $\Gamma_{+,\text{spin}}(t)$, $\Gamma_{-,\text{spin}}(t)$. A more detailed derivation of AME can be found in Appendix. A, while the charge and spin relaxation rates are computed in Appendix. B.
@@
$Q=1-2\overline{\big{|}\langle\tilde{\psi}_{\uparrow-}(t_{f})|{\tilde{\varrho}}(t_{f})|\tilde{\psi}_{\downarrow-}(t_{f})\rangle\big{|}},$ (14)
where the horizontal line represents classical averaging over distribution of $\delta t_{c}$, $\delta\epsilon$ and $\delta E_{Z,L/R}$.
@@ II.4 Parameters for GaAs and Si quantum dots
Charge relaxation time $\Gamma_{-}(2t_{c})$ $0.1\,\text{ns}^{-1}$ $10\,\text{ns}^{-1}$
@@
In this paper, we take into consideration two contrasting and experimentally relevant cases of isotopically purified Si and GaAs DQD devices. For both of them the relaxation rates $\Gamma_{-,\text{charge}}^{(s)}(t)$, $\Gamma_{-,\text{spin}}(t)$ are computed in Appendices B and C. In the polar GaAs the dominant role in the charge transfer is played by relatively strong coupling to piezoelectric phonons, which can improve charge transfer at low sweep rates [25]. However due to inevitable presence of nuclear spins, coherent transfer that takes longer than $T_{2}^{*}\approx 10$ns is impossible without additional control. For this reason we consider here GaAs device with a narrowed nuclear field [41]. In such case the effective dephasing time can be extended by a few orders of magnitude, i.e. $T_{2,\text{GaAs}}^{*}\approx 2\mu$s, at the cost of continuous estimation of the nuclear field [42]. We compare such a case against isotopically purified Si (with 800 ppm), with an order of magnitude longer $T_{2,\text{Si}}^{*}=20\mu$s [43].
To model experimental scenario we assume the detuning and tunnel coupling are affected by the charge noise of 1/f-type with a value of spectral density at $f=1$Hz, i.e. $S(1\text{Hz})\approx(0.5)^{2}\mu$eV2/Hz [34, 28], that translates to standard deviation of $\langle\delta\epsilon^{2}\rangle=5\mu$eV and $\langle\delta 2t_{c}\rangle=0.5\mu$eV. The assumed here, significantly weaker noise in tunnel coupling is consistent with a typical difference between the lever arm of plunger and barrier gates [44]. We add the second source of charge noise in form of Johonson noise from $R=50\Omega$ resistor. For spin-splitting noise we assume fluctuations in the dots $\delta E_{Z,L}$ and $\delta E_{Z,R}$ are independent and their standard deviation can be related to typically measured $T_{2}^{*}$, i.e. $\langle\delta E_{Z,L}^{2}\rangle=\langle\delta E_{Z,R}^{2}\rangle=\sqrt{2}/T_{2}^{*}$. We assume the $T_{2}^{*}$ is by the presence of the nuclei, or charge noise affecting the Zeeman splitting in presence of B field gradient [45, 46].
[…]
@@ III Sources of coherent transfer error
where $\Gamma_{-,\text{spin}}(t)$ is time-dependent spin-relaxation and pure dephasing contribution
@@
where $\langle\delta\varphi^{2}\rangle$ is the square of random phase $\delta\varphi$ averaged over realisations of low-frequency noise and random relaxation times. Each of them will be analyzed separately in the remainder of Sec. III and combined with the parameters corresponding to Si and GaAs devices in Sec. IV
@@ III.1 Charge transfer error $Q_{\text{charge}}$
The necessary condition for coherent communication is the high-fidelity charge transfer, i.e. deterministic motion of the electron between the dots. With sufficiently weak spin-orbit interaction and in the relevant regime of $2t_{c}>E_{Z}$, the spin components can be treated as uncoupled, and hence in the adiabatic regime (i.e. where $Q_{\text{LZ}}\ll 1$), the charge transfer error follows the result of [25]:
$Q_{\text{charge}}\approx\int_{0}^{t_{f}}\,\text{d}t\,\Gamma_{+}(\Omega[t])e^{-\int_{0}^{t}\Gamma_{+}(\Omega[t^{\prime}])+\Gamma_{-}(\Omega[t^{\prime}])\text{d}t^{\prime}}.$ (19)
The above equation can be derived from the AME from Eq. (11)), with the unitary part neglected, i.e. in the limit of $(2t_{c})^{2}\gg v$. At thermal equilibrium excitation and relaxation rates are related by the detailed balance condition $\Gamma_{+}(\Omega[t])=\Gamma_{-}(\Omega[t])e^{-\beta\Omega[t]}$ [49], where $\beta=1/k_{\text{B}}T$, so $\Gamma_{+}\!\ll\!\Gamma_{-}$ in most of the parameter regime of interest.
When coupling to environment is sufficiently weak, i.e. $\Gamma_{+}(2t_{c})t_{c}/v\ll 1$, one can consider only a single transition from ground to excited state, that is followed by a possible relaxation at finite detuning. This approach, termed Healed Excitation Approximation Limit (HEAL) in Ref. [25], results in a approximate formula:
@@
which shows that if healing factor ${\chi_{\text{HEAL}}=\int_{0}^{t_{f}}\Gamma_{-}(t^{\prime})\text{d}t^{\prime}\ll 1}$, the charge transfer error can be reduced by increasing $v$ since $Q\propto 1/v$, as larger $v$ reduces the time around avoided crossings, limiting excitations. On the other hand, slower sweeps, for which $\chi_{\text{HEAL}}>1$, enable correction of excitations through subsequent relaxation. This implies that $Q_{\text{charge}}$ can start decreasing below a certain $v$, as electrons in the excited state relax back to the ground state, reaching the target dot through inelastic tunneling, see Fig. 2). The subsequent sections will explore the extent to which such slow transfer of charge is accompanied by spin dephasing and spin relaxation.
@@ III.2 Spin relaxation error $Q_{\text{relax}}$
In this limit, the most relevant contribution comes from the mixing between |$\ket{\psi_{\downarrow+}(t)}$ and |$\ket{\psi_{\uparrow-}(t)}$ states around the avoided crossing. Since $E_{Z}>k_{\text{B}}T$ we concentrate on spin relaxation, i.e. inelastic transition from $|\psi_{\uparrow-}(t)\rangle$ to $|\psi_{\downarrow-}(t)\rangle$, which has dominant contribution from:
@@
The relaxation is upper bounded by its value at the avoided crossing. To compute probability of the spin-flip around avoided crossing $Q_{\text{relax}}$ from Eq. (17), we assume the relaxation rate is constant and non-negligible only around avoided crossing. Since the electron spends there a time period of the order of $2t_{c}/v$, we can estimate
@@
The figure indicates a significantly higher predicted probability of spin-relaxation in GaAs, compared to Si, attributed to larger intrinsic spin-orbit coupling (approximately an order of magnitude, $a=1\mu$eV) and a much faster relaxation rate at avoided crossings due to piezoelectric coupling. Even outside the manipulation region (red, with $b_{\perp}=0$), the resulting $Q_{\text{relax}}\geq 10^{-3}$ at $v=5\mu$eV/ns poses a serious threat to coherent transfer in GaAs. Conversely, the intrinsic spin-orbit interaction in Si (green) induces an effectively negligible error, $Q_{\text{relax}}<10^{-5}$ at $2t_{c}>40\mu$eV. Furthermore, Eq. (22) indicates that the intrinsic SOC contribution scales as $Q_{\text{relax}}\propto E_{Z}^{2}/t_{c}^{3}$, suggesting potential reduction with increased tunnel coupling or decreased magnetic field (Van Vleck cancellation).
However, if the dominant spin-orbit coupling contribution arises from the transverse gradient of the magnetic field $b_{\perp}$, with $b_{\perp}=1\mu$eV as an example corresponding to a gradient of $\sim 0.1$mT/nm and dots separation of $d\approx 100$nm. In such a case the scaling becomes $Q_{\text{relax}}\propto 1/t_{c}$. In GaAs (black), this contribution becomes significantly different from the intrinsic case only at large $2t_{c}$ (due to the lack of Van Vleck cancellation [50]). In Si (orange), it increases the relaxation rate by about two orders of magnitude, resulting in $Q_{\text{relax}}\approx 10^{-3}$ at $2t_{c}=40\,\mu$eV and $v=5\,\mu$eV/ns, potentially becoming a significant contribution to transfer error. It’s important to note that the error is expected to further increase as the tunnel coupling approaches $E_{Z}$, eventually becoming sensitive to random interference at $E_{Z}>2t_{c}$ due to additional avoided crossings.
Figure 3: Probability of spin-relaxation as a function of tunnel coupling for two different sweep rates $v=5,50\,\mu$eV/ns (solid and hollow dots). We compare numerical simulations of Adiabatic Master Equation (dots) against theoretical predictions from Eq. (17) (lines) for GaAs (red) and Si (green) devices, that are characterized by the relaxation rates from Appendix C. For each devices we compare contribution from intrinsic and synthetic spin-oribt coupling with the parameters given in Table 1. For the adiabatic master equation, we initialize the system in the equal superposition and assume effectively zero temperature, such that $\Gamma_{+}=0$. We confirm a good agreement between numerical and analytical methods in the relevant region of $2t_{c}>E_{Z}=30\,\mu$eV.

### v3

@@ II.3 Dynamics in presence of coupling to environment
Electric field noise leads to fluctuations of mean position of electron localized in each of the dots. In presence of finite gradient of $z$ component of the magnetic field, these lead to fluctuations in Zeeman splittings $E_{Z,L(R)}$ in both QDs [34, 35]. Similar fluctuations are caused by hyperfine interaction with nuclei in each of the dots [44, 45, 39]. We will include these processes later in the analysis, and now we will focus on processes that are activated by interdot tranfer of the electron caused by detuning sweep. These require the coupling of the environment to the charge states only:
@@
where the $\hat{V}_{\epsilon},\hat{V}_{t_{c}}$ operators acts on the environmental degrees of freedom and represent quantum fluctuations in detuning and tunnel coupling respectively. Such environment is typically assumed to be in thermal equilibrium at low, but finite electron temperature of $T\approx 100\,$mK such that $2t_{c}/k_{\text{B}}T\gtrapprox 1$ [1]. The high-frequency environmental noise causes inelastic transitions between the instantaneous eigenstates of Eq. (1) , which we model by Adiabatic Master Equation (AME) [46, 47, 48] of the form:
@@
where all of the operators and the states are written in the adiabatic frame, ${\tilde{O}}=\hat{A}(t)\hat{O}(t)\hat{A}^{\dagger}(t)$, |$\ket{\psi(t)}=\hat{A}(t)|\tilde{\psi}(t)\rangle$ i.e. in the basis associated with the instantaneous states of $\hat{H}_{0}(t)$ (see Sec.A.1 for more details). By definition, the transformation associated with an operator $\hat{A}(t)$ diagonalizes the Hamiltonian at every time instant:
@@
Firstly, we consider spin-diagonal transitions between the orbital states. After transformation the operators ${\tilde{\tau}}_{s\pm}=|\tilde{\psi}_{s\pm}(t)\rangle\langle\tilde{\psi}_{s\mp}(t)|$ are responsible for such transitions between the eigenstate of energy ${E_{s\pm}=\sigma_{s}E_{Z}/2\pm\Omega_{s}/2}$ and the eigenstate with energy ${E_{s\mp}=\sigma_{s}E_{Z}/2\mp\Omega_{s}/2}$, respectively. With this process we associate corresponding time-dependent rates $\Gamma_{\pm,\text{charge}}^{(s)}(t)$ of excitation/relaxation for spin-s. Secondly we add transitions between the eigenstates that are adiabatically linked to the opposite spins, i.e. ${\tilde{\sigma}}_{+}=|\tilde{\psi}_{\uparrow\pm}(t)\rangle\langle\tilde{\psi}_{\downarrow\pm}(t)|$, ${\tilde{\sigma}}_{-}=|\tilde{\psi}_{\downarrow\pm}(t)\rangle\langle\tilde{\psi}_{\uparrow\pm}(t)|$, with associated time-depedent transition rates $\Gamma_{+,\text{spin}}(t)$, $\Gamma_{-,\text{spin}}(t)$. A more detailed derivation of AME can be found in Appendix. A, while the charge and spin relaxation rates are computed in Appendix. B.
@@
$Q=1-2\Big{|}\overline{\langle\tilde{\psi}_{\uparrow-}(t_{f})|{\tilde{\varrho}}(t_{f})|\tilde{\psi}_{\downarrow-}(t_{f})\rangle}\Big{|},$ (14)
where the horizontal line inside the absolute value, represents classical averaging over distribution of $\delta t_{c}$, $\delta\epsilon$ and $\delta E_{Z,L/R}$.
@@ II.4 Parameters for GaAs and Si quantum dots
Charge relaxation rate $\Gamma_{-}(2t_{c})$ $0.1\,\text{ns}^{-1}$ $10\,\text{ns}^{-1}$
@@
In this paper, we take into consideration two contrasting and experimentally relevant cases of isotopically purified Si and GaAs DQD devices. They represent qualitatively different regimes in terms of charge relaxation rate (see analysis in [31]), but also in terms of interaction between the environment and spin degree of freedom analysed in this paper. For instance, due to relatively weaker spin-orbit interaction and the limited number of nuclear spins, isotopically purified silicon is considered a more promising candidate for coherent spin transfer.
For both of the material platforms the charge relaxation rates $\Gamma_{-,\text{charge}}^{(s)}(t)$, $\Gamma_{-,\text{spin}}(t)$ are computed in Appendix  C. In the polar GaAs the dominant role in the charge transfer is played by relatively strong coupling to piezoelectric phonons, which can improve charge transfer at low sweep rates [31]. However due to inevitable presence of nuclear spins, coherent transfer that takes longer than $T_{2}^{*}\approx 10$ns is impossible without additional control. For this reason we consider here GaAs device with a narrowed nuclear field [51, 52, 53]. In such a case the effective dephasing time can be extended by a few orders of magnitude, i.e. $T_{2,\text{GaAs}}^{*}\approx 2\mu$s, at the cost of continuous estimation of the nuclear field [54]. We compare such a case against isotopically purified Si (with 800 ppm), with an order of magnitude longer $T_{2,\text{Si}}^{*}=20\mu$s [55].
We assume the detuning and tunnel coupling are affected by the charge noise of 1/f-type with a value of spectral density at $f=1$Hz given by $S(1\text{Hz})\approx(0.5)^{2}\mu$eV2/Hz [42, 34], that translates to standard deviation of $\langle\delta\epsilon^{2}\rangle\approx 5\mu$eV. For tunnel coupling noise we assume significantly weaker amplitude $\langle\delta 2t_{c}\rangle\approx 0.5\mu$eV, which is consistent with a typical difference between the lever arm of plunger and barrier gates [56]. We add the second source of charge noise in form of Johnson noise from wiring with $50$ $\Omega$ impedance [31, 57]. For spin-splitting noise we assume fluctuations in the dots $\delta E_{Z,L}$ and $\delta E_{Z,R}$ are independent, and their standard deviation can be related to typically measured $T_{2}^{*}$, i.e. $\langle\delta E_{Z,L}^{2}\rangle=\langle\delta E_{Z,R}^{2}\rangle=\sqrt{2}/T_{2}^{*}$. We assume the $T_{2}^{*}$ is due the presence of the nuclei, or charge noise affecting the Zeeman splitting in presence of B field gradient [58, 59].
In both devices we use two values of tunnel coupling, $402t_{c}=$ 40 and $100$ $\mu$eV, and set $E_{Z}=30\mu$eV corresponding to $B\approx 0.25$T in Si, such that assumed condition $E_{Z}<2t_{c}$ is met. In Si we use two contrasting values of a difference in Zeeman splitting between the dots i.e. $\Delta E_{Z}=1,10^{-3}\mu$eV, where the first of them is more typical in presence of intentional gradient [60], for instance in the manipulation zone of the shuttling-based quantum computer architecture [61], while the latter can be associated with a typical difference in g-factors $\delta g/g\lesssim 10^{-3}$ [62]. For GaAs DQDs we use rather small $\Delta E_{Z}=0.1\mu$eV, but compare intrinsic spin-orbit coupling $a_{\text{im}}=1\mu$eV, against more optimistic $a_{\text{im}}=0.1\mu$eV. In Si we keep $a_{\text{im}}=0.1\mu$eV. To avoid interference we assume $a_{r}=0$. For both devices, we compare intrinsic spin orbit coupling against synthetic spin-orbit coupling $b_{\perp}$, produced by the longitudinal gradient of magnetic field, for instance in vicinity of manipulation region. We use value of $b_{\perp}\!=\!1\,\mu$eV, which in Si corresponds to the gradient of the order of 0.10.10.1\,0.1mT/nm and dots separation $d\approx 100\,$nm.
@@ III Sources of coherent transfer error
where $\Gamma_{-,\text{spin}}(t)$ is time-dependent spin relaxation and pure dephasing contribution
@@
where $\langle\delta\varphi^{2}\rangle$ is the square of random phase $\delta\varphi$ averaged over realisations of low-frequency noise and random time spent in the excited state. Each of them will be analyzed separately in the remainder of Sec. III and combined with the parameters corresponding to Si and GaAs devices in Sec. IV
@@ III.1 Charge transfer error $Q_{\text{charge}}$
Before considering spin degree of freedom, for completeness we start by a summary of results from Ref. [31], where we have analysed the necessary condition for coherent communication, i.e. the high-fidelity if the charge transfer. With sufficiently weak spin-orbit interaction and in the relevant regime of $2t_{c}>E_{Z}$, the spin components can be treated as uncoupled, and hence in the adiabatic regime (i.e. where $Q_{\text{LZ}}\ll 1$), the charge transfer error follows the result of [31]:
$Q_{\text{charge}}\approx\int_{0}^{t_{f}}\,\text{d}t\,\Gamma_{+}(\Omega[t])e^{-\int_{0}^{t}\big{(}\Gamma_{+}(\Omega[t^{\prime}])+\Gamma_{-}(\Omega[t^{\prime}])\big{)}\text{d}t^{\prime}}.$ (19)
The above equation can be derived from the AME from Eq. (11)), with the unitary part neglected, i.e. in the limit of $(2t_{c})^{2}\gg v$. It describes the occupation of the excited state obtained by integrating the excitation rate, which is normalised by the total number of transitions. For time-independent rates, the long-time limit would give the equilibrium occupation $Q_{\text{eq}}=\Gamma_{+}/(\Gamma_{-}+\Gamma_{+})$. However this prediction is expected to be heavily modified for time-dependent rates [63], for instance if $\Gamma_{+}$ is active only locally, around avoided crossing. This is often the case in thermal equilibrium, at which excitation and relaxation rates are related by the detailed balance condition $\Gamma_{+}(\Omega[t])=\Gamma_{-}(\Omega[t])e^{-\beta\Omega[t]}$ [64], where $\beta=1/k_{\text{B}}T$, so $\Gamma_{+}\!\ll\!\Gamma_{-}$ in most of the parameter regime of interest.
When coupling to environment is sufficiently weak, i.e. $\Gamma_{+}(2t_{c})t_{c}/v\ll 1$, one can consider only a single transition from ground to excited state, that is followed by a possible relaxation at finite detuning. This approach, termed Healed Excitation Approximation Limit (HEAL) in Ref. [31], resulted in a approximate formula:
@@
This result shows two ways of reducing incoherent charge transfer error. For relatively fast sweeps, i.e. for the healing factor ${\chi_{\text{HEAL}}=\int_{0}^{t_{f}}\Gamma_{-}(t^{\prime})\text{d}t^{\prime}\ll 1}$, the error can be reduced by increasing $v$ since $Q\propto 1/v$. This is because fast sweeps reduce the time around avoided crossings and hence limit the excitations. On the other hand, slower sweeps, for which $\chi_{\text{HEAL}}>1$, enable correction of excitations through subsequent relaxation. This implies that $Q_{\text{charge}}$ can start decreasing below a certain $v$, as electrons in the excited state relax back to the ground state, reaching the target dot through inelastic tunneling, see Fig. 2).
The subsequent sections will explore the extent to which such slow transfer of charge is accompanied by spin dephasing and spin relaxation.
@@ III.2 Spin relaxation error $Q_{\text{relax}}$
In this limit, the most relevant contribution comes from the mixing between |$\ket{\psi_{\downarrow+}(t)}$ and |$\ket{\psi_{\uparrow-}(t)}$ states around the avoided crossing. Since $E_{Z}>k_{\text{B}}T$ we concentrate on spin relaxation, i.e. inelastic transition between the states $|\psi_{\uparrow-}(t)\rangle$ and $|\psi_{\downarrow-}(t)\rangle$, which has dominant contribution from:
@@
The relaxation rate is upper bounded by its value at the avoided crossing. To compute probability of the spin-flip around avoided crossing $Q_{\text{relax}}$ from Eq. (17), we assume the relaxation rate is constant and non-negligible only around avoided crossing. Since the electron spends there a time period of the order of $2t_{c}/v$, we can estimate
@@
The figure indicates a significantly higher predicted probability of spin-relaxation in GaAs, compared to Si, attributed to larger intrinsic spin-orbit coupling (approximately by an order of magnitude, |$\absolutevalue{a}=1\mu$eV) and a much faster relaxation rate at avoided crossings due to piezoelectric coupling. Even outside the manipulation region (red, with $b_{\perp}=0$), the resulting $Q_{\text{relax}}\geq 10^{-3}$ at $v=5\mu$eV/ns poses a serious threat to coherent transfer in GaAs. Conversely, the intrinsic spin-orbit interaction in Si (green) induces an effectively negligible error, $Q_{\text{relax}}<10^{-5}$ at $2t_{c}>40\mu$eV. Furthermore, Eq. (22) indicates that the intrinsic SOC contribution scales as $Q_{\text{relax}}\propto E_{Z}^{2}/t_{c}^{3}$, suggesting potential reduction with increased tunnel coupling or decreased magnetic field (Van Vleck cancellation).
However, if the dominant spin-orbit coupling contribution arises from the transverse gradient of the magnetic field $b_{\perp}$, with $b_{\perp}=1\mu$eV as an example corresponding to a gradient of $\sim 0.1$mT/nm and dots separation of $d\approx 100$nm. In such a case the scaling becomes $Q_{\text{relax}}\propto 1/t_{c}$. In GaAs (black), this contribution becomes significantly different from the intrinsic case only at large $2t_{c}$ [65]. In Si (orange), it increases the relaxation rate by about two orders of magnitude, resulting in $Q_{\text{relax}}\approx 10^{-3}$ at $2t_{c}=40\,\mu$eV and $v=5\,\mu$eV/ns, potentially becoming a significant contribution to transfer error. It’s important to note that the error is expected to further increase as the tunnel coupling approaches $E_{Z}$, eventually being affected by interference effects [27], which are beyond the scope of this paper.
Figure 3: Probability of spin-relaxation as a function of tunnel coupling for two different sweep rates $v=5,50\,\mu$eV/ns (solid and hollow dots). We compare numerical simulations of Adiabatic Master Equation (dots) against theoretical predictions from Eq. (17) (lines) for GaAs (red) and Si (green) devices, that are characterized by the relaxation rates from Appendix C. For each device we compare contribution from intrinsic and synthetic spin-orbit coupling with the parameters given in Table 1. For the adiabatic master equation, we initialize the system in the excited spin state and assume effectively zero temperature, such that $\Gamma_{+}=0$. We confirm a good agreement between numerical and analytical methods in the relevant region of $2t_{c}>E_{Z}=30\,\mu$eV.